# Gravity versus astrophysics in black hole images and photon rings: equatorial emissions and spherically symmetric spacetimes

### v1

@@ 7 Astrophysical-geometrical degeneracy
If one knew exactly the emission process and the dynamics that govern the disc, deductions on the underlying geometry could be undoubtedly made through the electromagnetic signatures of the accretion onto the black hole. Nonetheless, bearing in mind our ignorance of the complex astrophysical mechanisms involved (Gralla 2021), the question is to know if, in spite of all the astrophysics uncertainties, there are still some observables that permit to distinguish between a Schwarzschild spacetime and a non-standard metric. In the two following paragraphs, we explore the feasibility of the previous statement using the two measured quantities introduced in Section 6, namely the width and the peak position of the observed intensity along the polar angles on the observer’s screen. We additionally separate the simplistic case in which the angle $\theta_{\mathrm{B}}$, between the direction of emission and the magnetic field, is fixed to be $90^{\circ}$ from the one in which it is computed numerically via equation (18.1), in order to examine the importance of including the variation in the $\theta_{\mathrm{B}}$ angle caused by the spacetime curvature. The conclusions made are very similar for both the observing frequencies considered in Table 3, see Appendix C for an explanation, and the plots presented correspond to images simulated at $\nu^{\mathrm{obs}}=345$ GHz.
@@ 7.1 Observable disentanglement of the peak positions
The direct and first-lensed peaks positions for all possible combinations of the emission parameters of Table 2 are confined in the different bands of Figures 7 and 8, each band corresponding to one of the spacetimes considered in Table 3: Schwarzschild in green, RZ with $a_{1}=0.1$ and $\epsilon=0$ in pink, RZ with $\epsilon=0.1$ and $a_{1}=0$ in violet. The three investigated velocities of the disc, specifically Keplerian, radial and mixed with a sub-Keplerian predominant part plus radial component, are explored separately in the three rows of the plots.
@@
Since the three bands are widely superposed in the first columns of Figures 7 and 8, we can state that astrophysical uncertainties dominate over small geometric deviations in the determination of the $n=0$ peaks and this regardless of the dynamics of the disc or the inclusion of $\theta_{\mathrm{B}}$. We remark that this is especially true for the Schwarzschild band, in green, and the $a_{1}|_{\,\epsilon=0}=0.1$ pink band which almost entirely cover each other because the event horizon, where all the emitted intensity profiles reach their maximum for a constant $\theta_{\mathrm{B}}$, see Section 6, is located at the same radial distance in these two cases. In other words, the other modifications induced by the non-standard metric, that is, on the geodesic motion, the redshift factor and the radius of the ISCO, see Section 6, are not sufficient to generate a significant observable difference on the $n=0$ peaks position and this is all the more true when varying only $b_{1}$, as illustrated in Appendix C. Even in the case of $\epsilon|_{a_{1}=0}=0.1$, whose event horizon’s radius is $r_{\mathrm{H}}\simeq 1.81M$, corresponding to a mean projected shift of about $0.84$ µas on the screen with respect to the Schwarzschild horizon, the measured positions of the peaks are highly degenerate between the green and violet bands. Higher parameters deviations lead, unsurprisingly, to more visible divergences, but even for the extreme positive values of equation (3.2), $a_{1}|_{\,\epsilon=0}=1$ and $\epsilon|_{a_{1}=0}=0.5$, the astrophysical-geometrical degeneracy is not broken for all the profiles, as shown in the first column of Figure C.7 of Appendix C.
@@
Figure 8: Positions of the $n=0$ (left panels) and $n=1$ (right panels) peaks, observed at 345 GHz and $i=163^{\circ}$, for all the possible emission configurations computed including the real value of the synchrotron pitch angle $\theta_{\mathrm{B}}$ for a vertical magnetic field in the case of a Schwarzschild black hole (green), a RZ black hole with $a_{1}|_{\,\epsilon=0}=0.1$ (pink) and a RZ black hole with $\epsilon\,|_{\,a_{1}=0}=0.1$ (violet). Each row corresponds to a different velocity: Keplerian, radial or mixed. Critical curves are represented as dashed lines.
@@
On the contrary, even being agnostic of the astrophysical emission, at least in the parameter space spanned in this paper, there is still the possibility to distinguish the strongly lensed observable features produced by different spacetime metrics. In other terms, as expected, the geometry of the spacetime impacts more dominantly the secondary images of the accretion disc. More precisely, for a non-standard black hole, the majority of the emission profiles with fixed $\theta_{\mathrm{B}}$ of a radially-infalling disc or of a disc with mixed velocity produce distinguishable peaks, see bottom right panels of Figure 7, and all of them become completely disentangled in the Keplerian case, see top right panel of Figure 7. Besides, if the real value of $\theta_{\mathrm{B}}$ is included, the degeneracy is always broken for all possible motions of the disc, as presented in the second column of Figure 8. This can be understood by the fact that subjacent spacetime plays an additional role in this second case, determining the lensing effects needed in the computation of $\theta_{\mathrm{B}}$. For a detailed explanation of the impact of $\theta_{\mathrm{B}}$ on the peak positions, also in the case of a radial magnetic field, see Appendix C. Again, as in the $n=0$ case, the distance from the Schwarzschild green band is more pronounced when the position of the event horizon is modified, as in the violet band [^25] We note that these violet peaks, contrary to the pink ones, would already be identified as a deviation from Schwarzschild when applying the purely geometric lensing band approach of Cárdenas-Avendaño & Held (2024), marginalising then over all conceivable emission profiles., and the contrasting conclusion vis-à-vis the $n=0$ study is mainly due to a modification of another key feature for the image formation, specifically the photon sphere around which $n=1$ light rays swirl before reaching infinity, see Section 6. This does not happen for deviations induced by the parameter $b_{1}$, see Appendix C. Here too, bigger parameters create more striking disentanglements, as shown in the second column of Figure C.7 of Appendix C. Also, a similar conclusion is reached when looking at higher order photon rings, see Figure C.1 in Appendix C.
@@
As a remark, we note that no curve of Figure 2 is retrieved exactly because the redshift factor deforms the emitted profiles differently for various values of the azimuthal angle, see Section 6, so the measured peaks along the polar angle on the screen are not simply the projections of circular rings as discussed in 2.3. Finally, in Appendices C.4 and C.5, we complement our analysis by presenting the impacts of an inner radius of the disc that does not match the radial position of the event horizon or of a drastically different inclination of the observer on our conclusions.
@@ 7.2 Tangle of the widths
The width is not an observable allowing us to disentangle the effects of the astrophysics and the geometry on the electromagnetic images of accretion discs around supermassive black holes. For small deviations they are completely degenerate and even when taking extreme allowed values of the metric parameters, the degeneracy is globally preserved, this for every image order and disc velocity profile, see two last columns of Figure C.7. This is not startling in view of the properties already highlighted in Section 6. In fact, the width is measured by taking the difference between the two radial distances at which the intensity is half of the maximum: a smaller one, affected by the gravitational redshift, being indeed sensitive to the variation of the metric, and a larger one being mostly prescribed by the falloff of the emission profile. To a certain extent, it is fair to say that the parametric deviations of the spacetime are drowned in the diversity of the considered radiation parameters.
@@ 8.1 Mass uncertainty
All the quantities of Figures 7 and 8 scale with the mass-to-distance ratio, $M/D$, which is expressed in angular units. From the data of the EHT 2017 campaign, the mass of M87* is measured to be $M=(6.5\pm 0.7)\times 10^{9}$ $\mathrm{M}_{\odot}$ under the assumption of GR and its distance is adopted to be $D=16.8\pm 0.8$ Mpc (The Event Horizon Telescope Collaboration 2019d). The precision on the mass reaches then about 11% whereas the one on the distance is close to 5%. Although the impact of the combined errors on $M$ and $D$ on our results should be examined, we note that the uncertainty on the mass is more than twice as important as that on the distance. Moreover, since the mass is the only parameter defining a Schwarzschild black hole, engendering the structure of its spacetime, we will assume, for the sake of simplicity, that the $D=16.9$ Mpc is perfectly known and that all the uncertainty on $M/D$ is concentrated on the mass. Thus, the question is to know whether a Schwarzschild black hole of lower mass, but still inside the EHT uncertainty range, could reproduce the $n=1$ characteristic peaks of a non-standard RZ black hole with the mass given in Table 1. To answer this, we put ourselves in the most optimistic scenario of Section 7, namely the one in which the disc has a mixed velocity and whose emissivity is given by expression (21), without further simplifications. Then, we superpose, in Figure 9, the bands of first-lensed peaks’ positions in the bottom right panel of Figure 8 to the ones produced by Schwarzschild black holes of lower masses: $M\in\{5.8,6.4,6.49\}$ $\mathrm{M}_{\odot}$, indicated by red ticks.
@@
We see that the present EHT precision on the value of the black hole mass is not sufficient to discriminate the underlying metric [^26] The current $M/D$ uncertainty also disables metric tests with photon rings for all alternative spherically symmetric BHs in Wielgus (2021). because small RZ geometric deviations, in dotted bands, can be imitated by a standard black hole with lower mass, contained well inside the uncertainty range. In other words, according to the EHT results, the peaks’ band linked to a Schwarzschild spacetime can appear everywhere between the green bands. However, improved mass measurements with error bars of the order of $0.1\times 10^{9}$ $\mathrm{M}_{\odot}$, corresponding to 2% on the red colour bar, could not replicate the violet peaks produced by a RZ metric with $\epsilon|_{a_{1}=0}=0.1$ and all the small spacetime deviations considered in this paper could be detected with a precision on the black hole mass value of the order of $0.01\times 10^{9}$ $\mathrm{M}_{\odot}$. To resume, small geometric deviations modifying the position of the event horizon could be detected by increasing the precision on the measured mass by about one order of magnitude, whereas one order of magnitude more is needed to disclose a generic small spacetime deviation. This conclusion is unvaried when considering the $n=2$ photon ring, as displayed in Appendix C.
@@
Conversely, assuming an infinite resolving power, the present EHT precision on the mass is sufficient to isolate the observables linked to the extreme $a_{1}|_{\epsilon=0}=1$ parametric deviation and the extreme positive deviation in $\epsilon$, $\epsilon|_{a_{1}=0}=0.5$, would become accessible too with an uncertainty on $M$ of about $0.4\times 10^{9}$ $\mathrm{M}_{\odot}$, that is, around 6% on the red colour bar. These results are illustrated by the starred bands in Figure 9. There exists then a one-to-one relationship between the value of the RZ parameters and the precision on $M/D$ needed to observe their effects.
@@
Figure 9: Positions of the $n=1$ peaks, observed at 345 GHz, for all the possible emission configurations in the case of a black hole of mass $M=6.5\times 10^{9}$ $\mathrm{M}_{\odot}$ described by a Schwarzschild metric (dark green), a RZ metric with $a_{1}|_{\,\epsilon=0}=0.1$ or $1$ (dotted pink or starred pink) and a RZ metric with $\epsilon\,|_{\,a_{1}=0}=0.1$ or $0.5$ (dotted violet or starred violet), and a Schwarzschild black hole of mass $M=5.8\times 10^{9}$ $\mathrm{M}_{\odot}$ (light green). The velocity of the disc is taken to be mixed and the synchrotron pitch angle is computed in the emissivity. The red colour bar shows the error bar from the mean black hole mass value, $M=6.5\times 10^{9}$ $\mathrm{M}_{\odot}$, while the grey colour bars indicate the value of the RZ parameters $a_{1}$ (pink) and $\epsilon$ (violet).
@@ 8.2 Relative measurements
To go around the stringent precision needed for the $M/D$-dependent measurements of slightly deviating spacetimes discussed in Section 8.1, for an image-plane analysis, one has to consider relative measurements linked for instance to the image asymmetry or to respective size of subsequent images. In order to construct such quantities, we introduce the diameter $d^{n}$ of an $n$-order image as the distance between two intensity peaks along a polar direction on the screen. We then define the fractional asymmetry $A^{n}$ of an $n$-order image as one minus the ratio between its smallest and biggest diameters and the respective size $R^{n}$ of successive $n$ and $n+1$ images as the ratio between their mean diameters:
@@
The asymmetry of the critical curve has been already investigated in Johnson et al. (2020) and proven to not exceed a few percent for the spinning black hole in M87, being zero percent for a non-rotating central object. We find that, under the assumed inclination, the asymmetry is around 1% and 3% for the primary image and it reaches at most 0.3% for the first photon ring. The latter behaviour is well explained by the convergence of photon rings to the circular critical curve, exacerbated as expected for higher order images as described in Appendix C. We note that, both in the $n=0$ and $n=1$ asymmetries, the astrophysical uncertainties are strongly dominating, preventing any test of GR via this measurement, even for the extreme geometric deviations presented in Appendix C. However, if precisions of about 0.2% could be reached on the asymmetry of the primary image, some constraints could be put on the velocity distribution of the disc, reducing then the astrophysical assumptions of our models. Also, as the asymmetry of n=1 is highly dependent on the inclination of the source [^27] And to a lesser extent on spin, as the critical curve (Gralla & Lupsasca 2020a)., see Appendix C for more details, it could be used as a direct measurement of it, complementary to the present one where the disc is assumed perpendicular to the direction of the jet in M87 (Mertens et al. 2016).
Similar conclusions are reached for the size ratios of subsequent images: these quantities are highly degenerate between different spacetimes, and even more for high inclinations. One possibility to try to reduce this degeneracy is to consider a $M/D$-dependent prior on the emission radius (Wielgus et al. 2020). Anyway, although not directly useful to perform tests of GR, relative photon-ring sizes could be used to constrain the value of the spin of a rotating black hole (Broderick et al. 2022) and reduce then again the astrophysical uncertainty.
To summarise, it is unlikely to test RZ deviations from Schwarzschild with relative measurements. The main impediment is the fact that under the spherical symmetry assumption, RZ deviations seem to act essentially as scaling modifications. Therefore, we cannot exclude more promising outputs in more exotic spacetimes with, for example, deformed lensing bands or non-proportional shifts of the event horizon and the photon shell.

### v2

@@ 7 Astrophysical-geometrical degeneracy
If one knew exactly the emission process and the dynamics that govern the disc, deductions on the underlying geometry could be undoubtedly made through the electromagnetic signatures of the accretion onto the black hole. Nonetheless, bearing in mind our ignorance of the complex astrophysical mechanisms involved (Gralla 2021), the question is to know if, in spite of all the astrophysics uncertainties, there are still some observables that permit one to distinguish between a Schwarzschild space-time and a non-standard metric. In the two following paragraphs, we explore the feasibility of the previous statement using the two measured quantities introduced in Section 6, namely the width and the peak position of the observed intensity along the polar angles on the observer’s screen. We additionally separate the simplistic case in which the angle between the direction of emission and the magnetic field, $\theta_{\mathrm{B}}$, is fixed to be $90^{\circ}$ from the one in which it is computed numerically via equation (18.1), in order to examine the importance of including the variation in the $\theta_{\mathrm{B}}$ angle caused by the space-time curvature. The conclusions made are very similar for both the observing frequencies considered in Table 3 (see Appendix C for an explanation), and the plots presented correspond to images simulated at $\nu^{\mathrm{obs}}=345$ GHz.
@@ 7.1 Observable disentanglement of the peak positions
The direct and first-lensed peaks positions for all possible combinations of the emission parameters of Table 2 were confined in the different bands of Figures 7 and 8, each band corresponding to one of the space-times considered in Table 3: Schwarzschild in green, RZ with $a_{1}=0.1$ and $\epsilon=0$ in pink, RZ with $\epsilon=0.1$ and $a_{1}=0$ in violet. The three investigated velocities of the disc, specifically Keplerian, radial, and mixed with a sub-Keplerian predominant part plus radial component, were explored separately in the three rows of the plots.
@@
Since the three bands are widely superposed in the first columns of Figures 7 and 8, we can state that astrophysical uncertainties dominate over small geometric deviations in the determination of the $n=0$ peaks and this regardless of the dynamics of the disc or the inclusion of $\theta_{\mathrm{B}}$. We remark that this is especially true for the Schwarzschild band, in green, and the $a_{1}|_{\,\epsilon=0}=0.1$ pink band which almost entirely cover each other because the event horizon, where all the emitted intensity profiles reach their maximum for a constant $\theta_{\mathrm{B}}$ (see Section 6), is located at the same radial distance in these two cases. In other words, the other modifications induced by the non-standard metric, that is, on the geodesic motion, the redshift factor, and the radius of the ISCO (see Section 6), are not sufficient to generate a significant observable difference on the $n=0$ peaks position and this is all the more true when varying only $b_{1}$, as illustrated in Appendix C. Even in the case of $\epsilon|_{a_{1}=0}=0.1$, whose event horizon’s radius is $r_{\mathrm{H}}\simeq 1.81M$, corresponding to a mean projected shift of about $0.84$ µas on the screen with respect to the Schwarzschild horizon, the measured positions of the peaks are highly degenerate between the green and violet bands. Higher parameters deviations lead, unsurprisingly, to more visible divergences, but even for the extreme positive values of equation (3.2), $a_{1}|_{\,\epsilon=0}=1$ and $\epsilon|_{a_{1}=0}=0.5$, the astrophysical-geometrical degeneracy is not broken for all the profiles, as shown in the first column of Figure C.7 of Appendix C.
@@
Figure 8: Positions of the $n=0$ (left panels) and $n=1$ (right panels) peaks, observed at 345 GHz and $i=163^{\circ}$, for all the possible emission configurations computed including the real value of the synchrotron pitch angle, $\theta_{\mathrm{B}}$, for a vertical magnetic field in the case of a Schwarzschild black hole (green), a RZ black hole with $a_{1}|_{\,\epsilon=0}=0.1$ (pink), and a RZ black hole with $\epsilon\,|_{\,a_{1}=0}=0.1$ (violet). Each row corresponds to a different velocity: Keplerian, radial or mixed. Critical curves are represented as dashed lines.
@@
On the contrary, even being agnostic of the astrophysical emission, at least in the parameter space spanned in this paper, there is still the possibility to distinguish the strongly lensed observable features produced by different space-time metrics. In other terms, as expected, the geometry of the space-time impacts more dominantly the secondary images of the accretion disc. More precisely, for a non-standard black hole, the majority of the emission profiles with fixed $\theta_{\mathrm{B}}$ of a radially infalling disc or of a disc with mixed velocity produce distinguishable peaks (see bottom right panels of Figure 7), and all of them become completely disentangled in the Keplerian case (see top right panel of Figure 7). Besides, if the real value of $\theta_{\mathrm{B}}$ was included, the degeneracy is always broken for all possible motions of the disc, as presented in the second column of Figure 8. This can be understood by the fact that subjacent space-time plays an additional role in this second case, determining the lensing effects needed in the computation of $\theta_{\mathrm{B}}$. A detailed explanation of the impact of $\theta_{\mathrm{B}}$ on the peak positions, also in the case of a radial magnetic field, is provided in Appendix C.3. Again, as in the $n=0$ case, the distance from the Schwarzschild green band is more pronounced when the position of the event horizon is modified, as in the violet band, [^25] We note that these violet peaks, contrary to the pink ones, would already be identified as a deviation from Schwarzschild when applying the purely geometric lensing band approach of Cárdenas-Avendaño & Held (2024), marginalising then over all conceivable emission profiles. and the contrasting conclusion vis-à-vis the $n=0$ study is mainly due to a modification of another key feature for the image formation, specifically the photon sphere around which $n=1$ light rays swirl before reaching infinity (see Section 6). This does not happen for deviations induced by the parameter $b_{1}$ (see Appendix C.7). Here too, bigger parameters create more striking disentanglements, as shown in the second column of Figure C.7 of Appendix C. Also, a similar conclusion was reached when looking at higher-order photon rings (see Figure C.1 in Appendix C).
@@
As a remark, we note that no curve of Figure 2 is retrieved exactly because the redshift factor deforms the emitted profiles differently for various values of the azimuthal angle (see Section 6) so the measured peaks along the polar angle on the screen are not simply the projections of circular rings as discussed in 2.3. Finally, in Appendices C.4 and C.5, we complement our analysis by presenting the impacts of an inner radius of the disc that does not match the radial position of the event horizon or of a drastically different inclination of the observer on our conclusions.
@@ 7.2 Tangle of the widths
The width is not an observable that allowed us to disentangle the effects of the astrophysics and the geometry on the electromagnetic images of accretion discs around supermassive black holes. For small deviations they are completely degenerate and even when taking extreme allowed values of the metric parameters, the degeneracy is globally preserved, this for every image order and disc velocity profile (see two last columns of Figure C.7). This is not startling in view of the properties already highlighted in Section 6. In fact, the width was measured by taking the difference between the two radial distances at which the intensity is half of the maximum: a smaller one, affected by the gravitational redshift, being indeed sensitive to the variation of the metric, and a larger one being mostly prescribed by the falloff of the emission profile. To a certain extent, it is fair to say that the parametric deviations of the space-time are drowned in the diversity of the considered radiation parameters.
@@ 8.1 Mass uncertainty
All the quantities of Figures 7 and 8 scale with the mass-to-distance ratio, $M/D$, which is expressed in angular units. From the data of the EHT 2017 campaign, the mass of M87* is measured to be $(6.5\pm 0.7)\times 10^{9} M\mathrm{M}_{\odot}$ under the assumption of general relativity and its distance is adopted to be $D=16.8\pm 0.8$ Mpc (The Event Horizon Telescope Collaboration 2019d). The precision on the mass thus reaches about 11%, whereas the one on the distance is close to 5%. Although the impact of the combined errors on $M$ and $D$ on our results should be examined, we note that the uncertainty on the mass is more than twice as important as that on the distance. Moreover, since the mass is the only parameter defining a Schwarzschild black hole, engendering the structure of its space-time, we assumed for the sake of simplicity that the $D=16.9$ Mpc is perfectly known and that all the uncertainty on $M/D$ is concentrated on the mass. Thus, the question is to know whether a Schwarzschild black hole of lower mass but still inside the EHT uncertainty range could reproduce the $n=1$ characteristic peaks of a non-standard RZ black hole with the mass given in Table 1. To answer this, we put ourselves in the most optimistic scenario of Section 7, namely the one in which the disc has a mixed velocity and whose emissivity is given by expression (21) without further simplifications. Then, we superposed in Figure 9 the bands of first-lensed peaks’ positions in the bottom-right panel of Figure 8 to the ones produced by Schwarzschild black holes of lower masses: $\in\{5.8,6.4,6.49\} M\mathrm{M}_{\odot}$ (indicated by red ticks).
@@
We observed that the present EHT precision on the value of the black hole mass is not sufficient to discriminate the underlying metric [^26] The current $M/D$ uncertainty also disables metric tests with photon rings for all alternative spherically symmetric black holes in Wielgus (2021). because small RZ geometric deviations, in dotted bands, can be imitated by a standard black hole with a lower mass contained well inside the uncertainty range. In other words, according to the EHT results, the peaks’ band linked to a Schwarzschild space-time can appear everywhere between the green bands. However, improved mass measurements with error bars of the order of $0.1\times 10^{9}$ $\mathrm{M}_{\odot}$, corresponding to 2% on the red colour bar, could not replicate the violet peaks produced by a RZ metric with $\epsilon|_{a_{1}=0}=0.1$, and all the small space-time deviations considered in this paper could be detected with a precision on the black hole mass value of the order of $0.01\times 10^{9}$ $\mathrm{M}_{\odot}$. To summarise, small geometric deviations modifying the position of the event horizon could be detected by increasing the precision on the measured mass by about one order of magnitude, whereas one order of magnitude more is needed to disclose a generic small space-time deviation. This conclusion does not change when considering the $n=2$ photon ring, as displayed in Appendix C.
@@
Conversely, assuming an infinite resolving power, the present EHT precision on the mass is sufficient to isolate the observables linked to the extreme $a_{1}|_{\epsilon=0}=1$ parametric deviation, and the extreme positive deviation in $\epsilon$, $\epsilon|_{a_{1}=0}=0.5$, would also become accessible with an uncertainty on $M$ of about $0.4\times 10^{9}$ $\mathrm{M}_{\odot}$, that is, around 6% on the red colour bar. These results are illustrated by the starred bands in Figure 9. Thus, there exists a one-to-one relationship between the value of the RZ parameters and the precision on $M/D$ needed to observe their effects.
@@
Figure 9: Positions of the $n=1$ peaks observed at 345 GHz for all the possible emission configurations in the case of a black hole of mass $6.5\times 10^{9} M\mathrm{M}_{\odot}$ described by a Schwarzschild metric (dark green), a RZ metric with $a_{1}|_{\,\epsilon=0}=0.1$ or 11 (dotted pink or starred pink), and a RZ metric with $\epsilon\,|_{\,a_{1}=0}=0.1$ or $0.5$ (dotted violet or starred violet) and a Schwarzschild black hole of mass $5.8\times 10^{9} M\mathrm{M}_{\odot}$ (light green). The velocity of the disc is taken to be mixed and the synchrotron pitch angle is computed in the emissivity. The red colour bar shows the error bar from the mean black hole mass value, $6.5\times 10^{9} M\mathrm{M}_{\odot}$, while the grey colour bars indicate the value of the RZ parameters $a_{1}$ (pink) and $\epsilon$ (violet).
@@ 8.2 Relative measurements
To go around the stringent precision needed for the $M/D$-dependent measurements of slightly deviating space-times discussed in Section 8.1, for an image-plane analysis, one has to consider relative measurements linked, for instance, to the image asymmetry or to respective size of subsequent images. In order to construct such quantities, we introduced the diameter of an $n$-order image, $d^{n}$, as the distance between two intensity peaks along a polar direction on the screen. We then defined the fractional asymmetry of an $n$-order image, $A^{n}$, as one minus the ratio between its smallest and biggest diameters and the respective size of successive $n$ and $n+1$ images, $R^{n}$, as the ratio between their mean diameters:
@@
The asymmetry of the critical curve has already been investigated in Johnson et al. (2020) and proven not to exceed a few percent for the spinning black hole in M87, being zero percent for a non-rotating central object. We find that under the assumed inclination, the asymmetry is around 1% and 3% for the primary image, and at most it reaches 0.3% for the first photon ring. The latter behaviour is well explained by the convergence of photon rings to the circular critical curve, and it is exacerbated, as expected, for higher-order images, as described in Appendix C. We note that both in the $n=0$ and $n=1$ asymmetries, the astrophysical uncertainties are strongly dominating, preventing any test of gravity via this measurement, even for the extreme geometric deviations presented in Appendix C. However, if precisions of about 0.2% could be reached on the asymmetry of the primary image, some constraints could be put on the velocity distribution of the disc, thus reducing the astrophysical assumptions of our models. Also, as the asymmetry of n=1 is highly dependent on the inclination of the source [^27] And to a lesser extent on spin, as the critical curve (Gralla & Lupsasca 2020a). (see Appendix C for more details), it could be used as a direct measurement of it, complementary to the present estimate where the disc is assumed to be perpendicular to the direction of the jet in M87 (Mertens et al. 2016).
We reached similar conclusions for the size ratios of subsequent images. These quantities are highly degenerate between different space-times and even more so for high inclinations. One possibility to try to reduce this degeneracy is to consider a $M/D$-dependent prior on the emission radius (Wielgus et al. 2020). Although not directly useful to perform tests of general relativity, relative photon-ring sizes could be used to constrain the value of the spin of a rotating black hole (Broderick et al. 2022) and then reduce again the astrophysical uncertainty.
To summarise, being able to test RZ deviations from Schwarzschild with relative measurements is unlikely. The main impediment is the fact that under the spherical symmetry assumption, RZ deviations seem to act essentially as scaling modifications. Therefore, we cannot exclude more promising outputs in more exotic space-times with, for example, deformed lensing bands or non-proportional shifts of the event horizon and the photon shell.